Frank is 4 years younger than Pam. The sum of their age is 80, write two linear equations to solve
How old are frank and Pam?

Answer by math_helper(2461)


Let P = Pam's age
P-4 = Frank's age
P+(P-4) = 80 —> P=42 —> F=38

Answer: Pam is 42, Frank is 38

Check: 42+38 = 80 (ok)